OpenStudy (anonymous):

y=4x^ +8x +3 What are x intercepts and Y intercept

OpenStudy (amistre64):

the last number is your y intercept: that +3. (0,3)

OpenStudy (amistre64):

you solve for the equation = 0 to get your x interecpts

OpenStudy (amistre64):

4x^2 +8x +3 = 0 4(3) = 12 1,12;2,6...2 and 6 work well. (x+6)(x+2) = 0 when either x+6 = 0 or when x+2 = 0

OpenStudy (amistre64):

your x intercepts are: (-6,0) (-3,0)

OpenStudy (anonymous):

what would be my vertex then?

OpenStudy (anonymous):

Vertex is at x = -b/2a which is -8/(2*4) = -1 so plug in x = -1 and you get y = 4(-1)^2 + 8(-1) + 3 which is -1 So your vertex is at (-1,-1)

OpenStudy (amistre64):

yeah, your vertex is simply halfway between your x intercepts

OpenStudy (amistre64):

hmmm...... i forgot to divide out my 4 again...... (x+6/4) (x+2/4) (x+3/2)(x+1/2) your x intercepts are (0,-3/2) (0,-1/2)

OpenStudy (amistre64):

you gotta let me know when I make a mistake...im not always gonna be around to catch these things lol

OpenStudy (amistre64):

like this blunder o'mine "(0,-3/2) (0,-1/2)" it should be (-3/2,0) (-1/2,0) brains going faster than my fingers can keep up :)